[Brahms-l] Shift report 20050512 00:00-08:00

From: <apache@brahms-web.brahms.bnl.gov>
Date: Thu May 12 2005 - 08:07:11 EDT
Shifter leader was rw
Shifters were 


1. Beam for 4h. The Beam was lost unexpectedly and the next filled
    store was lost in 30min again. 6 banch beam study was followed
    since then. MCR request us to change polarity because they think
    that all problem they have had since Tuesday may caused by our
    magnet. We changed the polarity on Tuesday. So now all 5 magnet
    polarity is A.

2. Data run 14592-14596.
   Pedestal runs 14597,14598

3. No dtector problem



_______________________________________________
Brahms-l mailing list
Brahms-l@lists.bnl.gov
http://lists.bnl.gov/mailman/listinfo/brahms-l
Received on Thu May 12 08:07:24 2005

This archive was generated by hypermail 2.1.8 : Thu May 12 2005 - 08:07:36 EDT